OpenStudy (anonymous):

Check it out Check it out Indicate in standard form the equation of the line passing through the given points. G(4, 6), H(1, 5)

OpenStudy (anonymous):

(y-y1) = m (x-x1) m = gradient = (y2-y1) / (x2-x1) = (5-6) / (1-4) = (-1) / (-3) = 1/3 (y-y1) = m (x-x1) where x1 and y1 is coordinate of point on the line. (y-4) = m (x-6) Hope that helps

OpenStudy (anonymous):

Yes Thanks